OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. (-4, -5) and (-4, 5)

OpenStudy (anonymous):

undefined because they have the same x-coordinates. this means the line that goes through them is vertical, which has an undefined slope

OpenStudy (anonymous):

and the work, it's simply m =(5-(-5))/(-4-(-4) = 10/0